Please help me solve this equation:
A triangle has a height of 2x and a base of 7-4x
What is the maximum area of the triangle?
(If somehow you can use quadratic formula to solve this that would be much appreciated)
Thank you so much!

Answer by Alan3354(69443)

A triangle has a height of 2x and a base of 7-4x
What is the maximum area of the triangle?
----------
Area = b*h/2
= 14x - 8x^2
dA/dx = 14 - 16x = 0
x = 7/8 units
------
max Area = (7 - 7/2)*(7/4)/2
max Area = 49/16 sq units